In the chemical reaction below, how many grams of HCl will be lrodiced by the reaction 2343.42g of CCl4? Show the solution.



CCl4 + 2HF -------> CCl2F2 + 2HCl

1
Expert's answer
2022-06-02T09:10:04-0400

Solution:

The molar mass of CCl4 is 153.82 g/mol

Therefore,

Moles of CCl4 = (2343.42 g CCl4) × (1 mol CCl4 / 153.82 g CCl4) = 15.2348 mol CCl4


Balanced chemical equation:

CCl4 + 2HF → CCl2F2 + 2HCl

According to stoichiometry:

1 mol of CCl4 produces 2 mol of HCl

Thus, 15.2348 mol of CCl4 produce:

(15.2348 mol CCl4) × (2 mol HCl / 1 mol CCl4) = 30.4696 mol HCl


The molar mass of HCl is 36.458 g/mol

Therefore,

Mass of HCl = (30.4696 mol HCl) × (36.458 g HCl / 1 mol HCl) = 1110.86 g HCl


Answer: 1110.86 grams of HCl will be prodiced by the reaction.